Question 6 18 6 Given A find A. 4 3
Answer #1

We have to find determinant of A

A=\begin{bmatrix} 8 & 6\\ 4 & 3 \end{bmatrix}

|A| = 8*3 - 4*6 = 24 - 24 = 0